OpenStudy (anonymous):

How do I solve 75y^3+100y^2-3y-4=0

OpenStudy (anonymous):

\[75y^{3}+100y^{2}-3y-4=0\]

jimthompson5910 (jim_thompson5910):

One way is to factor the left hand side

jimthompson5910 (jim_thompson5910):

we can factor by grouping to get 75y^3+100y^2-3y-4 (75y^3+100y^2)+(-3y-4) 25y^2(3y + 4) - 1(3y + 4) (25y^2 - 1)(3y + 4) (5y - 1)(5y + 1)(3y + 4) ------------------------------------------------------- So that means 75y^3+100y^2-3y-4 = 0 is the same as (5y - 1)(5y + 1)(3y + 4) = 0

jimthompson5910 (jim_thompson5910):

once you get to (5y - 1)(5y + 1)(3y + 4) = 0, you would use the zero product property

OpenStudy (anonymous):

Awesome. Thanks.

jimthompson5910 (jim_thompson5910):

sure thing